# MedRAX: Medical Reasoning Agent for Chest X-ray

### v1

@@ 1 Introduction
Chest X-rays (CXRs) have been widely used to make critical decisions in disease detection, diagnosis, and monitoring, comprising the largest proportion of over 4.2 billion diagnostic radiology procedures performed annually worldwide (United Nations Scientific Committee on the Effects of Atomic Radiation, 2022). However, the systematic evaluation of key anatomical structures places a significant time burden on radiologists, often requiring hours of careful analysis (Bahl et al., 2020).
@@
Foundation models (FMs), including large language models (LLMs) and large multimodal models (LMMs), have emerged as a promising solution to this fragmentation, enabling unified, scalable AI-driven image-text reasoning for medical tasks. OpenAI’s GPT-4 played a pivotal role in establishing the dominance of this approach with its unprecedented scale. Trained on an enormous volume of multimodal data, it has shown exceptional performance in medical understanding and reasoning without explicit training (Nori et al., 2023; Yan et al., 2023; Javan et al., 2024; Eriksen et al., 2024). LLaVA-Med (Li et al., 2024b), trained on 15 million biomedical figure-caption pairs, established new benchmarks in medical visual question answering (VQA), showcasing strong generalization in zero-shot image interpretation. CheXagent (Chen et al., 2024a) focused specifically on CXR analysis, achieving performance comparable to GPT-4 despite using significantly fewer parameters.
While FMs have significantly advanced the field, they face critical limitations that hinder their direct clinical application. LMMs frequently experience hallucinations and inconsistencies in their reasoning, particularly concerning for medical applications where accuracy is paramount. They also struggle with the complex multi-step reasoning required for diagnostic tasks, often failing to systematically evaluate all relevant anatomical structures or integrate findings across different regions of the image. Their end-to-end architecture, while elegant, lacks the transparency and specialization that comes from purpose-built medical AI tools. These limitations suggest the need for a more structured, tool-based approach that can combine the flexibility of foundation models with the reliability of clinical AI systems.
@@ 2.3 Evaluation Frameworks
Beyond general-purpose benchmarks, MedAgentBench (Jiang et al., 2025) assesses LLMs’ ability to retrieve patient data, interact with clinical tools, and execute structured decision-making in interactive healthcare environments. Results indicate that even the best-performing model, GPT-4o, achieves only 72% accuracy, with substantial performance variability across different medical tasks. These findings reinforce the need for domain-specific benchmarks that evaluate AI agents not just on general reasoning but on their ability to integrate into real-world clinical workflows.
@@ 3 MedRAX
Figure 1: Architecture of MedRAX. The framework implements a ReAct loop that processes user queries by integrating short-term memory (LangChain) and specialized medical tools for visual QA (CheXagent (Chen et al., 2024b), LLaVA-Med (Li et al., 2024b)), segmentation (MedSAM (Ma et al., 2024; Kirillov et al., 2023), ChestX-Det (Lian et al., 2021; Zhao et al., 2017)), grounding (Maira-2 (Bannur et al., 2024)), report generation (model trained on CheXpert Plus (Irvin et al., 2019; Chambon et al., 2024)), classification (TorchXRayVision (Cohen et al., 2022, 2020)), and image generation (RoentGen (Chambon et al., 2022)).
@@
     tool $=\text{SelectTool}(thoughts,T,M)$
     result $=\text{Execute}(tool,state)$
     $M=M\cup\{(thoughts,tool,result)\}$
     state $=\text{Observe}(state,result,M)$
@@ 3.1 LLM Driven Agent
MedRAX employs a LLM as the core to drive a ReAct (Reasoning and Acting) loop, which breaks down complex medical queries into sequential analytical steps (Yao et al., 2023). The system processes a user query through iterative cycles of (1) observation - analyzing the current state and query, (2) thought - determining required actions, and (3) action - executing relevant tools and integrating findings from previous steps to inform subsequent reasoning. Throughout this process, the system maintains a short-term memory of user interactions, tool outputs, and images to support multi-turn interactions. The reasoning loop continues until the system either generates a response or asks the user for additional input (Algorithm 1).
@@ 3.2 Flexible Tool Integration
    Models: MedSAM, a state-of-the-art biomedical segmentation model trained on 1,570,263 medical image-mask pairs, covering 10 imaging modalities and over 30 cancer types (Ma et al., 2024; Kirillov et al., 2023).
@@
The framework supports parallel execution of independent tools while providing flexible deployment configurations - tools can be quantized for efficiency and distributed across CPU or GPU resources. Figure 2 shows an example user interaction with MedRAX.
@@ 4 ChestAgentBench
Existing medical VQA benchmarks typically focus on simple, single-step reasoning tasks. In contrast, ChestAgentBench offers several distinctive advantages:
@@ 4.1 Dataset
From its chest imaging section, we curated 675 patient cases with associated chest X-rays and complete clinical documentation. These cases covered 53 unique areas of interest including lung, thorax, and mediastinum. Figure 3 provides an overview of the benchmark, showing (a) the creation pipeline, (b) patient gender distribution, (c) age distribution, and (d) most frequent anatomical areas of interest.
@@ 5 Experiments
Table 1: Model Performance on ChestAgentBench. Performance of five vision-language models (LLaVA-Med (Li et al., 2024b), CheXagent (Chen et al., 2024a), Llama-3.2-90B, GPT-4o, and MedRAX) compared across seven categories of our 2,500-question benchmark. MedRAX significantly outperforms both general-purpose models and specialized biomedical models across all categories.
@@
Table 2: Model Performance on CheXbench. Performance of five vision-language models (LLaVA-Med (Li et al., 2024b), CheXagent (Chen et al., 2024a), Llama-3.2-90B, GPT-4o, and MedRAX) compared on 238 Visual QA (Rad-Restruct and SLAKE) and 380 Image-Text Reasoning questions (OpenI). MedRAX excels in VQA while achieving the best overall performance.
@@ 5.1 Implementations
MedRAX uses GPT-4o as its backbone LLM and is deployed on a single NVIDIA RTX 6000 GPU, using the same configuration as described in Section 3. It integrates CheXagent (Chen et al., 2024b) and LLaVA-Med (Li et al., 2024b) for visual QA, Maira-2 for grounding (Bannur et al., 2024), a model trained on ChestX-Det for segmentation (Lian et al., 2021), TorchXRayVision for classification (Cohen et al., 2022), and a model trained on CheXpert Plus for report generation (Chambon et al., 2024).
MedRAX implements tool execution with structured JSON API calls, where the agent formulates precise requests with required arguments (e.g., image paths, text prompts) to call target tools. All baseline models are evaluated using their official implementations and recommended configurations.
Model responses are processed using regex to extract letter choices. For unclear responses, errors, or timeouts, we retry up to three times. Responses that remain invalid or do not choose a single choice are marked incorrect.
@@ 5.2 Experimental Setup
We evaluate MedRAX against four models: LLaVA-Med, a finetuned LLaVA-13B model for biomedical visual question answering ((Li et al., 2024b), CheXagent, a Vicuna-13B VLM trained for CXR interpretation (Chen et al., 2024b), along with GPT-4o and Llama-3.2-90B Vision as popular closed and open-source multimodal LLMs respectively.
We evaluate models on two complementary benchmarks:
@@ 5.4 Case Studies
This questions asks about diagnosing the predominant disease and comparing its severity across lungs. GPT-4o misinterprets the CXR as showing pneumonia with right lung predominance. MedRAX, through sequential tool application of report generation for disease identification and segmentation for lung opacity analysis, correctly determines left pneumothorax as the main finding. This demonstrates MedRAX’s ability to break down complex queries into targeted analytical steps.
@@ 6 Discussion
Our experiments demonstrate that MedRAX achieves state-of-the-art performance in complex CXR interpretation tasks, outperforming both general-purpose and specialized medical models. We discover valuable insights about structured tool orchestration in medical AI, suggesting that a hybrid approach—leveraging both large-scale reasoning capabilities and domain-specific expertise—offers superior performance over purely end-to-end models.
Task Decomposition. MedRAX demonstrates that a ReAct-based architecture dynamically composes complex reasoning chains while maintaining computational efficiency. The performance gap between MedRAX and end-to-end models suggests that explicit decomposition of reasoning steps provides advantages that scale alone cannot achieve. The process produces clear decision traces, enhancing transparency and interpretability, with implications beyond medical imaging for structured model-tool integration.
Generalists Versus Specialists. A key insight is the superior performance of general-purpose models (GPT-4o, Llama-3.2-90B) over specialized medical models (LLaVA-Med (Li et al., 2024b), CheXagent (Chen et al., 2024a)). This suggests that medical model specialization may sacrifice broader reasoning capabilities provided by large-scale pretraining. MedRAX bridges this gap by integrating domain-specific tools while maintaining generalist reasoning, demonstrating the benefits of hybrid architectures.
Limitations. While MedRAX excels in structured reasoning, it sometimes struggles with resolving contradictory tool outputs, particularly in fine-grained visual tasks when classification and segmentation tools provide conflicting interpretations of subtle patterns. Additionally, the system’s computational overhead from running multiple specialized tools can impact response times compared to end-to-end models. The framework also lacks robust uncertainty quantification mechanisms.
Future Work. Our initial observations suggest the importance of balanced tool utilization, where neither complete reliance on tools nor their complete absence produced optimal results. While formal analysis is needed, our findings indicate that prompting strategies encouraging critical evaluation of tool outputs may play a key role in system performance. This presents an interesting direction for understanding the interaction between LLM reasoning and tool utilization. Future work would consider applying reinforcement learning to boost reasoning ability, as verified in DeepSeek-R1 (Guo et al., 2025). Additionally, comprehensive clinical validation will be crucial for establishing MedRAX’s practical utility in real-world settings.
@@ 7 Conclusion
MedRAX establishes a new benchmark in AI-driven CXR interpretation by integrating structured tool orchestration with large-scale reasoning. Our evaluation on ChestAgentBench demonstrates its superiority over both general-purpose and domain-specific models, reinforcing the advantages of explicit stepwise reasoning in medical AI. These findings highlight the potential of combining foundation models with specialized tools, a principle that could be applied to broader domains in healthcare and beyond. Future work should focus on optimizing tool selection, uncertainty-aware reasoning, and expanding MedRAX’s capabilities to multimodal medical imaging for greater clinical impact.
8 Acknowledgements

### v2

@@ 1 Introduction
Chest X-rays have been widely used to make critical decisions in disease detection, diagnosis, and monitoring, comprising the largest proportion of over 4.2 billion diagnostic radiology procedures performed annually worldwide (United Nations Scientific Committee on the Effects of Atomic Radiation, 2022). However, the systematic evaluation of key anatomical structures places a significant time burden on radiologists, requiring hours of careful analysis (Bahl et al., 2020; Adibi et al., 2025).
@@
Foundation models (FMs), including large language models (LLMs) and large multimodal models, are promising solutions to this fragmentation, enabling unified, scalable AI-driven medical image-text reasoning. OpenAI’s GPT-4 established dominance of this approach with unprecedented scale. Trained on enormous multimodal data, it demonstrated exceptional medical understanding and reasoning without explicit training (Nori et al., 2023; Yan et al., 2023; Javan et al., 2024; Eriksen et al., 2024; Baghbanzadeh et al., 2025). LLaVA-Med (Li et al., 2024b), trained on 15 million biomedical figure-caption pairs, established new benchmarks in medical visual question answering (VQA) with strong zero-shot image interpretation. CheXagent (Chen et al., 2024a) focused on CXR analysis, achieving GPT-4-level performance with significantly fewer parameters.
While FMs have advanced the field, they face critical limitations hindering direct clinical application. LMMs experience hallucinations and inconsistencies in reasoning, particularly concerning where medical accuracy is paramount. They struggle with complex multi-step reasoning required for diagnostic tasks, failing to systematically evaluate all relevant anatomical structures or integrate findings across different image regions. Their end-to-end architecture lacks the transparency and specialization of purpose-built medical AI tools. These limitations suggest a more structured, tool-based approach combining foundation model flexibility with clinical AI system reliability.
@@ 2.2 Medical Agents
Alongside these agentic frameworks, specific large vision-language models like RaDialog (Pellegrini et al., 2025) have been developed for radiology report generation and conversational assistance. Furthermore, models such as M4CXR (Park et al., 2024) are exploring the multi-task potentials of multi-modal LLMs for chest X-ray interpretation.
@@ 2.3 Evaluation Frameworks
Beyond general-purpose benchmarks, MedAgentBench (Jiang et al., 2025) assesses LLMs’ ability to retrieve patient data, interact with clinical tools, and execute structured decision-making in interactive healthcare environments. Results indicate that even the best-performing model, GPT-4o, achieves only 72% accuracy, with substantial performance variability across different medical tasks.
AgentClinic (Schmidgall et al., 2024) offers a comprehensive multimodal agent benchmark to evaluate AI in simulated clinical environments, emphasizing interactive dialogue and active data collection processes. These findings reinforce the critical need for domain-specific benchmarks and highlight that while progress has been made, current medical agents still face significant challenges in tool integration, efficiency, and flexibility.
@@ 3 MedRAX
Figure 1: Architecture of MedRAX. The framework implements a ReAct loop that processes user queries by integrating short-term memory (LangChain) and specialized medical tools for visual QA (CheXagent (Chen et al., 2024b), LLaVA-Med (Li et al., 2024b)), segmentation (MedSAM (Ma et al., 2024; Kirillov et al., 2023; Ma et al., 2025), ChestX-Det (Lian et al., 2021; Zhao et al., 2017)), grounding (Maira-2 (Bannur et al., 2024)), report generation (model trained on CheXpert Plus (Irvin et al., 2019; Chambon et al., 2024)), classification (TorchXRayVision (Cohen et al., 2022, 2020)), and image generation (RoentGen (Chambon et al., 2022)).
@@
     tools $=\text{SelectTools}(thoughts,T,M)$
     results $=\text{ExecuteParallel}(tools,state)$
     $M=M\cup\{(thoughts,tools,results)\}$
     state $=\text{Observe}(state,results,M)$
@@ 3.1 LLM Driven Agent
MedRAX employs a LLM as the core to drive a ReAct (Reasoning and Acting) loop, which breaks down complex medical queries into sequential analytical steps (Yao et al., 2023). The system processes a user query through iterative cycles of (1) observation - analyzing the current state and query, (2) thought - determining required actions, and (3) action - executing relevant tools and integrating findings from previous steps to inform subsequent reasoning. Throughout this process, the system maintains a short-term memory of user interactions, tool outputs, and images to support multi-turn interactions. The reasoning loop continues until the system either generates a response or asks the user for additional input (Algorithm 1). Further details of the core methodology are provided in Appendix A.
@@ 3.2 Flexible Tool Integration
    Models: MedSAM, a state-of-the-art biomedical segmentation model trained on 1,570,263 medical image-mask pairs, covering 10 imaging modalities and over 30 cancer types (Ma et al., 2024, 2025).
@@
The framework supports parallel execution of independent tools and flexible deployment configurations - tools can be quantized for efficiency and distributed across CPU or GPU. Figure 2 shows an example user interaction with MedRAX.
@@ 4 ChestAgentBench
While MedRAX offers a powerful framework for complex CXR interpretation, existing medical VQA benchmarks typically focus on simple, single-step reasoning tasks, making them insufficient for evaluating its full capabilities. In contrast, ChestAgentBench offers several distinctive advantages:
@@ 4.1 Dataset
From its chest imaging section, we curated 675 patient cases with associated chest X-rays and complete clinical documentation. These cases cover 53 unique areas of interest including lung, thorax, and mediastinum. Figure 3 provides an overview of the benchmark, showing (a) the creation pipeline, (b) patient gender distribution, (c) age distribution, and (d) most frequent anatomical areas of interest. The dataset encompasses diverse clinical settings and a broad spectrum of pathologies, with detailed breakdowns provided in Appendix B.
@@ 4.2 Benchmark Creation
We performed automated quality verification using GPT-4o to evaluate each question for: (1) structural consistency (six-choice format with one correct answer), (2) explicit grounding in the clinical and radiological context, and (3) clear verifiability from the original Eurorad case material. Questions failing these criteria were automatically excluded.
@@ 5 Experiments
Table 1: Model Performance on ChestAgentBench. Accuracy (%) of five vision-language models (LLaVA-Med (Li et al., 2024b), CheXagent (Chen et al., 2024a), Llama-3.2-90B, GPT-4o, and MedRAX) compared across seven categories of our 2,500-question benchmark. MedRAX significantly outperforms both general-purpose models and specialized biomedical models across all categories.
@@
Table 2: Model Performance on CheXbench. Accuracy (%) of five vision-language models (LLaVA-Med (Li et al., 2024b), CheXagent (Chen et al., 2024a), Llama-3.2-90B, GPT-4o, and MedRAX) compared on 238 Visual QA (Rad-Restruct and SLAKE) and 380 Image-Text Reasoning questions (OpenI). MedRAX excels in VQA while achieving the best overall performance.
@@ 5.1 Implementations
MedRAX uses GPT-4o as its backbone LLM, and we deploy it on a single NVIDIA RTX 6000 GPU using the same configuration as described in Section 3. It integrates CheXagent (Chen et al., 2024b) and LLaVA-Med (Li et al., 2024b) for visual QA, Maira-2 for grounding (Bannur et al., 2024), a model trained on ChestX-Det for segmentation (Lian et al., 2021), TorchXRayVision for classification (Cohen et al., 2022), and a model trained on CheXpert Plus for report generation (Chambon et al., 2024).
MedRAX implements tool execution with structured JSON API calls, where the agent formulates precise requests with required arguments (e.g., image paths, text prompts) to call target tools. We evaluate all baseline models using their official implementations and recommended configurations.
We process model responses using regex to extract letter choices. For unclear responses, errors, or timeouts, we retry up to three times. Responses that remain invalid or do not choose a single choice are marked incorrect.
@@ 5.2 Experimental Setup
We evaluate MedRAX against four core models: LLaVA-Med, a finetuned LLaVA-13B model for biomedical visual question answering (Li et al., 2024b), CheXagent, a Vicuna-13B VLM trained for CXR interpretation (Chen et al., 2024b), along with GPT-4o and Llama-3.2-90B Vision as popular closed and open-source multimodal LLMs respectively. Additionally, we compare against baseline models from recent literature including RadFM (Wu et al., 2023), a generalist foundation model for radiology, MAIRA-1 (Hyland et al., 2024), a specialized multimodal model for radiology report generation, LLaVA-Rad (Zambrano Chaves et al., 2024), a radiology-adapted version of LLaVA, and Med-PaLM M 84B (Tu et al., 2023), a large-scale biomedical foundation model, with performance values obtained from M4CXR paper (Park et al., 2024).
We evaluate models on four complementary benchmarks:
@@
(3) MIMIC-CXR Radiology Report Generation, which evaluates single-image chest X-ray findings generation of 3,858 images from the MIMIC-CXR test set. This benchmark assesses clinical accuracy of generated reports using micro-averaged F1 scores (mF1-14, mF1-5) and macro-averaged F1 scores (MF1-14, MF1-5) for 14 medical observation labels and 5 key findings (cardiomegaly, edema, consolidation, atelectasis, pleural effusion), respectively.
(4) SLAKE VQA, which evaluates medical visual question answering using 114 chest X-ray test samples with close-ended questions in English, filtered from the original SLAKE test set of 2,094 samples. Performance is measured by accuracy (exact matches) and recall (proportion of ground truth words present in generated responses).
For more information on preparation and evaluation on benchmarks 3 and 4, see M4CXR paper (Park et al., 2024).
Table 3: Single-image performance on MIMIC-CXR test set. Clinical accuracy (%) of generated radiology findings, evaluated using CheXbert F1 scores for 14 medical observations. mF1-14/mF1-5: micro-averaged F1 scores for all 14/5 key labels (cardiomegaly, edema, consolidation, atelectasis, pleural effusion); MF1-14/MF1-5: macro-averaged F1 scores for all 14/5 key labels. MedRAX is compared with CheXagent (Chen et al., 2024b), MAIRA-1 (Hyland et al., 2024), LLaVA-Rad (Zambrano Chaves et al., 2024), Med-PaLM M 84B (Tu et al., 2023), and M4CXR (Park et al., 2024). Performance values for baseline models obtained from Park et al. (Park et al., 2024).
Model mF1-14 mF1-5 MF1-14 MF1-5
Med-PaLM M 84B 53.6 57.9 39.8 51.6
CheXagent 39.3 41.2 24.7 34.5
MAIRA-1 55.7 56.0 38.6 47.7
LLaVA-Rad 57.3 57.4 39.5 47.7
M4CXR 60.6 61.8 40.0 49.5
MedRAX 79.1 64.9 34.2 48.2
Table 4: Medical VQA performance on SLAKE benchmark. Accuracy (%) and recall (%) of vision-language models on 114 chest X-ray visual question answering samples with close-ended English questions. MedRAX achieves state-of-the-art performance compared to RadFM (Wu et al., 2023), CheXagent (Chen et al., 2024b), and M4CXR (Park et al., 2024). Performance values for baseline models obtained from Park et al. (Park et al., 2024).
Model Accuracy Recall
RadFM 68.4 69.7
CheXagent 71.1 73.2
M4CXR 85.1 86.0
MedRAX 90.35 91.23
@@ 5.3 Quantitative Analysis
MIMIC-CXR Radiology Report Generation. Shown in Table 3, MedRAX achieves the highest micro-averaged F1 scores with mF1-14 of 79.1% and mF1-5 of 64.9%, substantially outperforming the next-best M4CXR (60.6%, 61.8%) and other baselines. However, MedRAX shows lower macro-averaged performance, with M4CXR achieving the highest MF1-14 of 40.0% and Med-PaLM M 84B leading MF1-5 at 51.6%. This suggests MedRAX’s stronger performance on more prevalent conditions.
SLAKE VQA. Shown in Table 4, MedRAX demonstrates clear superiority in medical visual question answering, achieving 90.35% accuracy and 91.23% recall. This represents significant improvements over M4CXR (85.1% accuracy, 86.0% recall) and CheXagent (71.1% accuracy, 73.2% recall). The consistent performance gains across both metrics indicate MedRAX’s tool-based approach effectively handles diverse medical VQA tasks while maintaining high precision in response generation.
@@ 5.4 Case Studies
This questions asks about diagnosing and comparing disease severity across lungs. GPT-4o misinterprets the CXR as showing pneumonia with right lung predominance. MedRAX, through sequential tool application of report generation for disease identification and segmentation for lung opacity analysis, correctly determines left pneumothorax as the main finding. This demonstrates MedRAX’s ability to break down complex queries into targeted analytical steps.
@@ 6 Discussion
MedRAX achieves state-of-the-art performance in complex CXR interpretation tasks, outperforming both general-purpose and specialized medical models. We discover valuable insights about structured tool use in medical AI, suggesting that a hybrid approach—leveraging both large-scale reasoning capabilities and domain-specific expertise—offers superior performance over purely end-to-end models.
Task Decomposition. MedRAX shows that the ReAct loop dynamically composes complex reasoning chains while maintaining computational efficiency. The performance gap suggests that explicit decomposition provides advantages that scale alone cannot achieve. The process produces clear decision traces, enhancing transparency with implications beyond medical imaging for model-tool integration.
Generalists Versus Specialists. A key insight is the superior performance of general-purpose models (GPT-4o, Llama-3.2-90B) over specialized medical models (LLaVA-Med (Li et al., 2024b), CheXagent (Chen et al., 2024a)). This suggests medical specialization may sacrifice broader reasoning capabilities from large-scale pretraining. MedRAX bridges this gap by integrating domain-specific tools while maintaining generalist reasoning.
Limitations. While MedRAX excels in structured reasoning, it sometimes struggles with resolving contradictory tool outputs, particularly in fine-grained visual tasks when classification and segmentation tools provide conflicting interpretations. Additionally, the system’s computational overhead from running multiple specialized tools can impact response times compared to end-to-end models. The framework also lacks robust uncertainty quantification mechanisms.
Future Work. Key areas for advancement include formally analyzing optimal tool utilization balance, as neither full reliance nor complete absence yielded best results. This involves exploring trade-offs between accuracy gains and computational costs, speed, error propagation, and spurious correlations in pretrained models. Systematic investigation of prompting techniques for critical tool evaluation is needed to optimize LLM-tool interactions. Reinforcement learning approaches could enhance reasoning capabilities and reduce hallucinations, following advances in models like DeepSeek-R1 (Guo et al., 2025). Comprehensive clinical validation remains essential for establishing real-world utility.
@@ 7 Conclusion
MedRAX establishes a new benchmark in AI-driven CXR interpretation by integrating structured tool orchestration with large-scale reasoning. Our evaluation demonstrates its superiority over both general-purpose and domain-specific models, reinforcing the advantages of explicit stepwise reasoning in medical AI. These findings highlight the potential of combining foundation models with specialized tools, a principle applicable to broader healthcare domains. Future work should focus on optimizing tool selection, uncertainty-aware reasoning, and expanding capabilities to multimodal medical imaging for greater clinical impact.
Impact Statement
MedRAX aims to improve chest X-ray interpretation, enhancing diagnostic accuracy and efficiency as an AI co-pilot designed to augment, not replace, clinical expertise. We acknowledge AI risks, including model bias, hallucinations, and data privacy, and emphasize the need for robust validation before any clinical deployment to ensure patient safety.
Our development utilized the de-identified Eurorad database, and its use, along with ChestAgentBench creation, strictly adheres to the CC BY-NC-SA 4.0 license; ChestAgentBench will be released under the same terms. To address data privacy, especially for datasets like MIMIC-CXR requiring stricter handling, we employed the Azure OpenAI Service configured to opt-out of data logging, following Physionet recommendations.
MedRAX also supports local LLM deployment for maximum data control. Our public code release will include guidelines for these privacy-preserving configurations, underscoring our commitment to responsible AI integration in healthcare.
Acknowledgements